APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday May 18, 2003 (MJD 52778) ---=== OBSERVING TEAM ===--- Daniel Long (APO; swing) Mike Harvanek (APO) Steph Snedden (APO) Carlos Gonzalez (by phone) Wendell Jordan (by phone) Jon Brinkmann (by phone) ---=== OBSERVING PLAN ===--- Shakedown and, if possible, science. ---=== OBSERVING SUMMARY ===--- Most of shake completed but spectro on-sky testing/science was cut short due to a problem with sp1 power supply. Engineering staff was notified of the problem and they will work on it today. On-sky imager testing and collimation not done but will be done tomorrow. Problems with sp1 power supply required cycling power every few hours so that the secondary dewars would fill. Southwest corner of the enclosure making a clunking noise (see log). East door making a sctratching noise (see log). Steph set up the Dimm and ran it for the last half of the night. She will continue to run it (inside its enclosure) during the day. ---=== OBSERVING LOG ===--- Afternoon Report: ================= Summary of yesterday: Tape drives cleaned. Spectro monthly checkout was completed, and the data looked reasonable. biasGen and biasCheck with the imager done. This afternoon the spectrographs were refocused. Below are the mean offset values from collimate before and after pistoning by 3000. Before After b1: .21 .01 b2: .40 .16 r1: -.35 -.11 r2: -.32 -.09 Final Piston Position sp1 sp2 motor A 8513 16666 motor B 8529 16636 motor C 8514 16647 I had to power cycle the guider as it was not talking to the guider Mac ( ppk err 86 ). This fixed the problem. I needed to home the primary because D and E were off by about .2mm. Secondary was fine as it was. Note that the D MiG is off by more than the nominal .05mm. It has been off by an ever increasing amount with successive dark runs and this month went beyond the normally accepted error. This is almost certainly not due to a change in the position of the primary, rather it is the quill engraving a line in the puck and slowly changing the length the quill extends. If it is found that the collimation check with the imager on sky looks good, I would recommend generating a new set of nominal values for the primary MiGs. Running plotMcpFidcuials showed a change from last months positions which indicate approximately 6 arcseconds of error accumulated in azimuth in 180 degrees due to scale error. A 90 degree sweep in altitude gave a similar scale error ( 3 arcseconds ). Neither of these scale errors of ~.03 arcseconds per degree were deemed large enough to make a change to the fiducial tables. Night: ====== While waiting for twilight to end, all changes listed in the changeLog after version v1_67 were verified. Plate 1223 (cartridge 3): With all offsets cleared, an FK5 star appeared on the edge of fiber 10 on the gotoField. Precalibs failed and after trying several things (dropping and grabbing instruments, logging out ports, rebooting mechanicals) we discovered sp1 had a bad power supply. We cycled power but the supply did not come back up. After talking to Carlos G., Wendell J. and Jon B. on the phone, we learned that the spare supply was in the process of being rebuilt. Since the moon would be rising soon and it was a shake night anyway, we all agreed that the problem could be dealt with in the morning. Jon reminded us that the secondary dewars would not fill so we decided we would cycle power on the sp1 supply every few hours to fill them. We cycled power to make sure this would work and it did. It also brought back the dead power supply. Unfortunately, the supply went belly up again shortly thereafter. Steph thinks the bad supply is moth-related. Since moon rise was near, we put the enclosure back on to obtain LED data. While putting the enclosure on, a small clanking noise was heard from the southwest corner. It began about 10ft from the west end of the rails, occurred about once per wheel revolution and lasted until it reached the west end. Visual inspection turned up nothing. Also, as the east door of the enclosure was being closed, it made a scraping or scatching sound that began about a foot from the bottom and lasted until it reached the bottom. The east door was rolled up a few feet and then back down and the noise was repeatable. We then obtained LED data with the camera in the closed doghouse with the enclosure on, closed and darkened. The t-bars were unlatched. One frame from each ccd was stored in /data/photo/cmr/52778. The flat vectors were examined by eye (checkLED and plotFlatvec) and u3 is still as ugly as ever. However, since we don't have nominal vectors for comparison, we can't be sure that things haven't changed. The power supply for sp1 was power cycled at 0655Z and the dewars filled for 60-70 secs. This time the top 3 supplies had red warning lights and were dead (rather than just a single supply). A bias in the doghouse with the t-bars unlatched was done. One frame from each ccd was stored in /data/photo/cmr/52778. checkBias gave 1 failure and 4 warnings: im> checkBias 3945 26 /data/photo/cmr/52778 CAMERA im: FAILURE on ftelnet: connection error in ftelnet: connection refused {11 WARN: } {11 FAIL: leftBiasVector rightBiasVector } {22 WARN: rightBiasVector} {22 FAIL: } {51 WARN: leftBiasVector } {51 FAIL: } {53 WARN: leftBiasVector } {53 FAIL: } {56 WARN: rightBiasVector} {56 FAIL: } The mosaic file was displayed and nothing looked unusual. The bias vectors were plotted for each ccd that gave a warning or failure in checkBias but nothing looked that far out of the ordinary. Since no changes were made to the fiducial tables this month and an FK5 star was found in fiber 10 immediately on the gotoField (with previous offsets cleared), we decided not to do a new pointing model tonight. If we find our pointing to be off tomorrow night we will do a new model then. Steph played around with the Dimm tonight. The pointing was not very accurate but she was finally able to get it set up on Polaris again. It will remain on Polaris through the day (although its dome will be closed) and we will check on it/play with it again tomorrow night. The power supply for sp1 was power cycled at 0925Z and the dewars filled for 50-60 secs. This time only a single supply (-15V ?) had a red warning light and was dead. The primary D mig nominal value was changed from -9.1130 (old nominal value) to -9.0560 (tonight's current value) as per Dan's suggestion (see above). We ran endNight and shipped tapes due to the LED calibration data. The sp1 power supply was power cycled at 1115Z and the dewars filled for about 60 secs. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 3944 05:24Z 05:40Z 10 O -64.00 -64.00 38 calibration 3945 06:56Z 07:10Z 100 O 49.86 53.42 38 bias ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): n UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL6393 - 6398, 6405 Stays: JL6399 - 6404, 6406 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 01:49Z 3 0 1.00000 0 1257 0 121 30.0 18.0 12 288 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:44Z 70 24 19 263 (W) 692 - Cirrus 00:17Z 68 22 20 255 (WSW) 767 - Cirrus 00:49Z 67 20 24 251 (WSW) 908 - Cirrus 01:37Z 65 16 17 272 (W) 1175 - pcl 02:08Z 64 15 11 268 (W) 1068 - " 02:38Z 63 15 13 289 (WNW) 1139 - " 03:08Z 62 16 13 277 (W) 1098 - " 03:38Z 62 15 14 270 (W) 1194 - " 04:10Z 61 15 10 275 (W) 1234 - " 04:43Z 61 16 7 268 (W) 1281 - " 05:15Z 61 17 11 280 (W) 1180 - " 05:47Z 60 17 11 280 (W) 1265 - " 06:19Z 60 19 10 275 (W) 1176 - " 06:51Z 59 19 17 268 (W) 1258 - " 07:24Z 57 20 10 316 (NW) 1403 - " 07:56Z 57 22 13 338 (NNW) 1352 - clear 08:28Z 57 22 10 14 (NNE) 1494 - clear 09:00Z 57 22 6 3 (N) 1507 - clear 09:32Z 56 21 8 6 (N) 1582 - clear 10:05Z 56 21 2 345 (NNW) 1700 - pcl 10:37Z 56 21 8 0 (N) 1734 - pcl 11:09Z 56 21 16 340 (NNW) 1680 - pcl ---=== TELESCOPE STATUS ===--- 01:30Z Enclosure doors opened, fans and blowers on, louvers opened. 02:10Z Enclosure off. 05:15Z Enclosure on (gathering LED data). 10:35Z Fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cartridge 3 (plate 1223) mounted. Counterweights at 290. Spectro auto-fill connected. Imager dewar 11:19 144 lb Spectro dewar 11:19 132 lb No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_113_1 Watcher: v2_22_0 MCP: v5_18_0 TPM: tpm_v2_29_0 AstroDa: v14_47 TCC: 2.6.8 sdssProcedures: v1_69 tccMon: v2_0 SoS: v4_9_13 hoggPT: v1_6_8 plate-mapper: v4_2_0 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0950 0.0790 Axial B 0.8020 0.7970 Axial C 0.8210 0.8240 Trans D -9.0560 -9.1130 Lateral E 1.8490 1.8870 Lateral F 0.0000 1.4300 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5377. -3699. 923. -208. 31535. 30647. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 19.7 Alt.: 30.001649 MIGS TONIGHT NOMINAL Axial A 1.5090 1.5170 Axial B 1.1510 1.1560 Axial C 1.0950 1.0990 Trans D -0.6990 -0.7070 GALILS Commanded: 1593136. 1530805. 1539800. -3400. -7900. Actual: 1619636. 1555273. 1576326. -3383. -7925. SETMIR VALUES SecDesOrient: 1257.00 0.00 -20.00 0.00 130.98 SecOrient: 1256.38 0.74 -21.40 -1.40 130.39 ---=== PROBLEMS IN DETAIL ===---